# ExAMPC: the Data-Driven Explainable and Approximate NMPC with Physical Insights

### v1

@@ VI Conclusion
This work introduces ExAMPC, an explainable and approximate NMPC and monitor framework for NMPC-controlled autonomous systems operating under model mismatch and environmental uncertainties. One aim of the work is to assist users without deep technical expertise to easily comprehend and operate an NMPC and its behavior. We propose the embedding of time-series through a Legendre-Spline encoding for dimensionality reduction, to approximate and explain the open-loop primal solutions of the NMPC as state and control trajectories, through a physics-inspired loss, enhancing the continuous-time safety satisfaction by 93%. Additionally, by combining SHAP and Symbolic Regression, ExAMPC provides an explainable performance monitor to uncover the physical insights affecting performance indicators such as closed-loop cost value and the impact of vehicle yaw rate on the computation time. Future work could leverage these explainability results for targeted data generation in edge cases using DTs, and integrate SR-derived analytical KPI models directly into the NMPC optimization.

### v2

@@ VI Conclusion
This work introduces ExAMPC, an explainable and approximate NMPC and monitor framework for NMPC-controlled autonomous systems operating under model mismatch and environmental uncertainties. One aim of the work is to assist users without deep technical expertise to easily comprehend and operate an NMPC and its behavior. We propose the embedding of time-series through a Legendre-Spline encoding for dimensionality reduction, to approximate and explain the open-loop primal solutions of the NMPC as state and control trajectories, through a physics-inspired loss, enhancing the continuous-time safety satisfaction by 93%, and achieving close to no constraints violation. Additionally, by combining SHAP and Symbolic Regression, ExAMPC provides an explainable performance monitor to uncover the physical insights affecting performance indicators such as closed-loop cost and predicted energy value and the impact of measurements such as vehicle yaw rate and tracking errors on the computation time. Future work could leverage these explainability results for targeted data generation in edge cases using DTs, and integrate SR-derived analytical KPI models directly into the NMPC optimization.